# G2SDF: Surface Reconstruction from Explicit Gaussians with Implicit SDFs

### v1

@@ C.1 Results on DTU and Mip-NeRF 360
In Table 5, we present the results reported in the Neuralangelo publication [50] on the DTU dataset. It is important to note that the quantitative results from the Neuralangelo paper have been flagged by other studies and discussions on the official GitHub repository as non-reproducibleRefer to this GitHub issue for details..
@@
C.2 Mesh Extraction with Different Rasterizers
@@
In Figure 9, we present meshes extracted using various Gaussian rasterizers, showcasing the versatility and adaptability of our method. These results demonstrate that our approach can be seamlessly integrated into existing Gaussian methods, no matter it’s 2D Gaussians [1] or 3D Gaussians [2], enhancing their performance without requiring significant modifications. Furthermore, our method consistently delivers high-quality reconstructions with fine details, highlighting its effectiveness in capturing intricate scene structures such as transparent and reflection areas.
Appendix D Limitation Discussion
@@
The pseudo depth maps exhibit high uncertainty in the background regions and inconsistency across multiple views, which limits our method’s ability to reconstruct distant background areas, such as the sky. However, since foreground objects are the primary focus in reconstruction tasks, our method demonstrates superior performance in these regions. To address this limitation, we plan to introduce an uncertainty-based weighting mechanism for the geometry regularization.
@@
Figure 9: Mesh Extraction with Different Gaussian Rasterizer.

### v2

@@ C.1 Results on DTU and Mip-NeRF 360
In Table 7, we present the results reported in the Neuralangelo publication [21] on the DTU dataset. It is important to note that the quantitative results from the Neuralangelo paper have been flagged by other studies and discussions on the official GitHub repository as non-reproducible [^2] Refer to this GitHub issue for details..
@@
Appendix D More Ablation
Figure 7: Mesh Extraction with Different Gaussian Rasterizer.
D.1 Ablation on Different Rasterizers
@@
In Figure 7, we present reconstruction results using different Gaussian rasterizers, showcasing the versatility and adaptability of our method. These results demonstrate that our approach can be seamlessly integrated into existing Gaussian methods, no matter it’s 2D Gaussians [12] or 3D Gaussians [48], enhancing their performance without requiring significant modifications. Furthermore, our method consistently delivers high-quality reconstructions with fine details, highlighting its effectiveness in capturing intricate scene structures such as transparent and reflection areas.
D.2 Ablations on Normalization Function.
[…]
D.3 Ablations on SDF to Opacity Function.
We compare results using 3DGSR’s [24] SDF-to-Opacity function (denoted as “w/ 3DGSR SDF2O” in Table 6) with our proposed function (Eq. 5). Our method achieves a higher f-score, offering a simpler, more efficient solution and the ability to reach full opacity of 1, unlike 3DGSR’s capped opacity of 0.25. While 3DGSR targets object-level reconstruction, this limitation may impede its effectiveness in unbounded outdoor scenes.
Method w/ MipNeRF NF w/ 3DGSR SDF2O Ours
F-score 0.52 0.24 0.56
Table 6: Ablation Study on Normalization and SDF to Opacity. We show the F-score of reconstructed mesh on Barn.
D.4 Ablations on Geometry Cues.
Our analysis reveals significant artifacts in Depth Anything v2’s estimated depth maps due to inherent scale uncertainty, despite global alignment attempts. As shown in Figure 8, we compare our alignment method “Ours” with DN Splatter’s approach [34] “DN Align”, demonstrating our method’s superior ability to reduce depth errors between rendered and aligned pseudo depth. Interestingly, despite the inaccuracies in DN Splatter’s aligned pseudo depth, the reconstruction F1 score remains comparable to our method. This finding suggests that while these limitations underscore the unreliability of pseudo depth maps for precise reconstruction, necessitating our model’s strong ability to compensate for misalignments, the primary improvement stems from our novel Gaussian-SDF collaboration pipeline rather than geometric cues alone.
Figure 8: Ablations on Geometry Cues. We show the ground truth color image, rendered depth, aligned pseudo depth and depth error (in log) on Barn. DN Align represents our pipeline with DN Splatter’s [34] alignment.
D.5 Ablations on Pruning Strategy.
Our method links the opacity value of each Gaussian to it’s SDF value with our SDF-to-Opacity function. While this maintains computational equivalence to conventional opacity-based pruning, it provides a physically-grounded criterion.
D.6 Ablations on Multi-Resolution Regularization.
Our Multi-Resolution training strategy employs a coarse-to-fine approach, initially utilizing fewer Gaussians at lower resolutions to establish robust scene geometry before progressively refining details. Figure 10 quantitatively illustrates the controlled growth of Gaussian populations throughout training. While the quantitative results remain the same since the evaluation only focus on foreground objects, the qualitative results shows significant improvement. Figure 9 demonstrates the effectiveness of our approach, showing that multi-resolution regularization enables reliable reconstruction of objects with limited views, whereas standard training struggles in such scenarios.
Figure 9: Ablation Study on Multi-Resolution Regularization. We show a comparison of ground truth color images alongside our rendered color, depth, and normal outputs at various training iterations on Barn. The results demonstrate that our multi-resolution regularization enables more efficient and complete scene geometry reconstruction, achieving superior convergence compared to standard training approaches.
Figure 10: Total Gaussian Population Growth on Barn (Training Iterations vs. Gaussian Count). Our proposed multi-resolution regularization achieves comparable reconstruction and rendering quality with significantly fewer Gaussians than standard training approaches, demonstrating improved computational efficiency without compromising output quality.
D.7 Ablations on Different Meshing Methods.
Figure 11 illustrates the limitations of depth fusion-based mesh extraction in unbounded scenes, revealing significant distortion caused by resolution constraints and MipNeRF’s [3] normalization artifacts. Comparative results in Figure 12 demonstrate our method’s advantages across different meshing techniques, showing that our Gaussian-guided surface reconstruction achieves both superior efficiency and enhanced detail preservation compared to existing approaches.
Figure 11: Comparison on Depth Fusion and Ours on Barn.
Figure 12: Ablation Study on Different Meshing Methods. While Marching Cubes struggles with resolution limitations in unbounded scene reconstruction, resulting in low-quality meshes with missing details, our method consistently produces high-quality meshes with enhanced geometric fidelity and preserved fine structures.
Appendix E Limitation Discussion
@@
The pseudo depth maps exhibit high uncertainty in the far distance regions and inconsistency across multiple views, which limits our method’s ability to reconstruct distant background areas, such as the sky. However, since foreground objects are the primary focus in reconstruction tasks, our method demonstrates superior performance in these regions. To address this limitation, we plan to introduce an uncertainty-based weighting mechanism for the geometry regularization.